OpenStudy (anonymous):

anyone in the 9th grade or above

OpenStudy (anonymous):

I am

OpenStudy (anonymous):

oka what is the hypothesis and conclusion of 5z+2=12 when z=2

OpenStudy (anonymous):

hypothesis and concluesion that is science

OpenStudy (anonymous):

are u looking for how to do the problem?

OpenStudy (australopithecus):

Replace z in the equation with 2

OpenStudy (anonymous):

no its in math..for example:if 2x=12 then x=6

OpenStudy (anonymous):

ikt

OpenStudy (anonymous):

and no

OpenStudy (australopithecus):

I know for a fact that there are people on this site that are in graduate programs to answer your initial question. To answer your second question 5z+2=12 when z=2 you get 5(2) + 2 = 12

OpenStudy (anonymous):

don't worry u guys don't understand i go to an honors school and ima freshman..algebra 1